The proportion of eligible voters in the next election who will vote for the ANC is assumed to be 0.55 in Gauteng. What is the probability that in a random of 500 Gauteng voters less than 0.49 say they will vote for the ANC?
1
Expert's answer
2021-07-21T12:58:03-0400

If both of np and nq are greater than or equal to 10, then the sampling distribution of sample proportions will be approximately normal with mean p and variance "\frac{pq}{n}" .

"\hat{p}~ N(P, \frac{pq}{n})"

p = population proportion,

q = 1 - p

n = sample size

p = 0.55

q = (1 - 0.55) = 0.45

n = 500

np = 275, which is greater than 10.

nq = 225, which is greater than 10.

Hence sampling distribution of sample proportions will be approximately normal.

We have to find P("\hat{p} < 0.49" ).

"Z = \frac{\hat{p}-p}{\sqrt{\frac{pq}{n}}} \\

P(\hat{p}<0.49) = P( \frac{\hat{p}-p}{\sqrt{\frac{pq}{n}}} < \frac{0.49-p}{\sqrt{ \frac{pq}{n} }}) \\

= P(Z< \frac{0.49-0.55}{\sqrt{ \frac{0.55 \times 0.45}{500} }}) \\

= P(Z< -2.696) \\

= 0.035"